OpenStudy (anonymous):

differentiate = sinx/(1-2cosx)

OpenStudy (anonymous):

\[\frac{d(\frac{\sin x}{1 - 2\cos x})}{dx} = \frac{\cos x(1 - 2\cos x) - \sin x(2\sin x)}{(1 - 2\cos x)}\]

OpenStudy (anonymous):

assume cosx=t & then proceed